# pre-trained model AST

### v1

@@
PickleBall: Secure Deserialization of Pickle-based Machine Learning Models
Andreas D. Kellas Columbia UniversityNew YorkNYUSA Neophytos Christou Brown UniversityProvidenceRIUSA Wenxin Jiang Purdue UniversityWest LafayetteINUSA Socket Penghui Li Columbia UniversityNew YorkNYUSA Laurent Simon GoogleMountain ViewCAUSA Yaniv David Technion – Israel Institute of TechnologyHaifaIsrael Vasileios P. Kemerlis Brown UniversityProvidenceRIUSA James C. Davis Purdue UniversityWest LafayetteINUSA  and  Junfeng Yang Columbia UniversityNew YorkNYUSA
@@ Abstract.
Machine learning model repositories such as the Hugging Face Model Hub facilitate model exchanges. However, bad actors can deliver malware through compromised models. Existing defenses such as safer model formats, restrictive (but inflexible) loading policies, and model scanners have shortcomings: 44.9% of popular models on Hugging Face still use the insecure pickle format, 15% of these cannot be loaded by restrictive loading policies, and model scanners have both false positives and false negatives. Pickle remains the de facto standard for model exchange, and the ML community lacks a tool that offers transparent safe loading.
We present PickleBall to help machine learning engineers load pickle-based models safely. PickleBall statically analyzes the source code of a given machine learning library and computes a custom policy that specifies a safe load-time behavior for benign models. PickleBall then dynamically enforces the policy during load time as a drop-in replacement for the pickle module. PickleBall generates policies that correctly load 79.8% of benign pickle-based models in our dataset, while rejecting all (100%) malicious examples in our dataset. In comparison, evaluated model scanners fail to identify known malicious models, and the state-of-art loader loads 22% fewer benign models than PickleBall. PickleBall removes the threat of arbitrary function invocation from malicious pickle-based models, raising the bar for attackers to depend on code reuse techniques.
@@
conference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwanccs: Security and privacy Software and application security
@@ 1. Introduction
Open-source and open-weight models enable the AI ecosystem (Jiang et al., 2022; Davis et al., 2023; Wang et al., 2022a; Han et al., 2021). They enable machine learning engineers to exchange pre-trained models rather than training from scratch (Dubey et al., 2024), facilitating direct use or further fine-tuning (Church et al., 2021). Open model repositories like Hugging Face now host millions of pre-trained models for many tasks (Jiang et al., 2023a, 2022). These model hubs are accessed directly as well as through corporate mirrors (Zhao et al., 2024), with billions of downloads per month.
Unfortunately, as with traditional software supply-chain attacks, bad actors can distribute malicious models. The most common strategy for achieving remote code execution is by tampering with the model deserialization process. Several model serialization formats, such as TorchScript (PyTorch, 2024b), H5/HDF5 (The HDF Group, 2025), and the Python pickle module (Python, 2025b), permit executable metadata or callbacks during model deserialization. Attackers can craft malicious serialized models to execute code, such as os.system(), on victim systems during model loading (Casey et al., 2024; Cohen, 2024; Wickens and Bonner, 2024; Dark Reading Staff, 2024). Researchers have found malicious pickle models on Hugging Face whose payloads include system fingerprinting, credential theft, and reverse shells (Cohen, 2024; Wickens and Bonner, 2024; Zhao et al., 2024), with one study discovering a $5{\times}$ increase in the rate of malicious models uploaded to Hugging Face year-over-year (Zhao et al., 2024). This led to alternative safe model formats like SafeTensors (Hugging Face, 2024b), and restricted loading APIs like the PyTorch weights-only unpickler (PyTorch, 2024c); we study their adoption in the Hugging Face ecosystem (§3) and find that nevertheless, insecure formats are still prevalent.
In this work, we propose a novel approach to secure pickle model deserialization, which we focus on for three reasons. First, pickle is a popular exchange format for models: repositories with pickle models are downloaded over 2.1 billion times per month from the Hugging Face model hub (§3.1). Second, pickle is the most expressive format and thus presents challenges to secure; it encodes models as opcodes executed by the pickle virtual machine (the Pickle Machine(of Bits, 2024; Huang et al., 2022) during deserialization, and permits invocations of arbitrary Python classes and functions (callables). Finally, pickle is a known security target for attackers, with almost all malicious models discovered on Hugging Face using the pickle format.
@@
To address these limitations, we present PickleBall, a two-part system for securing the exchange of pickle-based models. Our insight is that we can analyze library code to determine the expected behaviors of benign models produced by the library, and enforce tailored model loading policies. PickleBall statically analyzes the library code to learn (1) all classes used in the library’s models and their transitive attributes, and (2) all functions for restoring objects of these classes from serialized bytes. Then, PickleBall’s model-loading component enforces the inferred policies.
We evaluated PickleBall and the state-of-the-art approaches on a dataset of 336 models. Our dataset is meant represent the kinds of models that PickleBall must handle, including malicious models from attackers presenting as benign models used for inference tasks. We used 252 benign models sourced from Hugging Face, and 84 real and synthetic malicious models. PickleBall loads 79.8% of benign models for inference tasks and prevents all malicious models from executing their payloads. PickleBall adds a runtime overhead of $\sim$2.62% to safely load a model. Compared to other approaches, PickleBall achieved favorable precision and recall.
@@
    The design and implementation of PickleBall, a framework for securely loading pickle models. PickleBall tailors loading policies to models, and enforces these policies lazily, for secure, efficient, and robust model loading.
@@
    A novel dataset of 336 benign and malicious pickle-based models for evaluating pickle security efforts. It has 252 benign models collected from Hugging Face, and 84 malicious models. [^1] Code and data will be published on acceptance and provided for artifact review.
@@ 2.1.1. The Model Supply Chain
Training ML models from scratch requires massive compute and data (Patterson et al., 2021; Dubey et al., 2024), so engineers and companies save resources by reusing machine-learning models trained by others (pre-trained models) (Jiang et al., 2022; Davis et al., 2023; Wang et al., 2022a; Han et al., 2021). Open model repositories like Hugging Face host over 1.8M models (Hugging Face, 2025b) for many tasks (Jiang et al., 2023a, 2022).
A supply chain of pre-trained models has grown from model reuse, which comes with risks similar to those in the traditional software supply chain (Ohm et al., 2020; Jiang et al., 2022). Bad actors apply techniques familiar in traditional software security, such as typosquatting (Neupane et al., 2023; Liu et al., 2022; Jiang et al., 2024, 2025) and code injection (Wang et al., 2022b), as well as ML-specific ones like data manipulation (Gu et al., 2019). Model hubs like Hugging Face try to detect malicious models using both traditional code scanners like ClamAV (Jiang et al., 2022) and domain-specific pickle scanners (AI, 2024a; Hugging Face, 2024a) due to the proliferation of malicious pickle models (Casey et al., 2024; Zhao et al., 2024).
@@ 2.1.2. Model Serialization Formats
Model savers and loaders must agree on the serialization format; there are various formats available, each with its own tradeoffs in terms of security, flexibility, and performance. 222 A convenient table showing these tradeoffs is provided in the SafeTensors repository README: https://github.com/huggingface/safetensors. Python is the primary language for using ML models, and its native serialization module, pickle (Python, 2025b), provides a convenient and flexible interface for saving objects; pickle proliferated for being easy to use and is used by popular libraries like PyTorch (PyTorch, 2024a). Hugging Face released the SafeTensors format in September 2022 as an alternative that prioritizes security (Hugging Face, 2024b). The GGUF format, released August 2023, is optimized for fast model loading and inference tasks, especially for large language models (GGML, 2025; Hugging Face, 2025a). Other formats may be selected for library coupling (e.g., TensorFlow SavedModel), interoperability (e.g., ONNX), or intermediate tradeoffs between security, flexibility, and performance.
The security of a format depends on the expressivity of the operations needed to deserialize a model. The SafeTensors format requires very few different operations to load a model, because it effectively only encodes model weight values, and is considered a secure format after independent security audits (Trail of Bits, 2023; Patry, Nicolas and Biderman, Stella, 2023). Some formats, like TensorFlow SavedModel and ONNX, are known to permit operations that could be abused in specific settings (Zhu et al., 2025; Travers, 2021), but with no observed real-world attacks. Pickle is an extremely expressive format that permits nearly arbitrary operations during deserialization, and numerous malicious pickle models are discovered on Hugging Face (Zhao et al., 2024; Casey et al., 2024; Zanki, 2025; Cohen, 2024; Wickens and Bonner, 2024).
We focus our efforts on pickle models due to their insecure format (explained in §2.2) and their continued popularity (measured in §3).
@@ 2.2. Pickle Serialization and Risks
import pickle
def read_weights_to_tensor(filename: str) -> Tensor:
# Read a file containing weights and
# return a Tensor object.
class Tensor(object):
def __reduce__(self):
return (read_weights_to_tensor, self.filename)
class Model(object):
def __init__(self, weights: library.Tensor):
self.weights = weights
def save(self, filename):
pickle.dump(filename, self)
@classmethod
def load(path)
return pickle.load(path)
@@
Pickle Deserialization Attacks: Pickle’s opcodes allow a pickle program to import and invoke arbitrary Python callables during unpickling. In particular, an attacker can make use of pickle’s callable importing and callable invoking opcodes to achieve arbitrary code execution—for example by executing an arbitrary shell command by invoking the os.system function. The dangers of deserializing arbitrary pickle programs have been publicized for over a decade (ColdwaterQ, 2022; of Bits, 2024; Elhage, 2023; Python, 2025b).
Protecting Pickle Models: Two existing approaches help engineers load pickle models more securely:
@@
    Model scanners identify malicious models using denylists of unsafe callables. As with many denylist approaches, model scanners are useful for identifying recognizable malicious content, but are bypassed by malicious models that avoid denied callables, or that indirectly invoke callables (Tumarkin, 2024). [^3] We demonstrate this in Appendix B by creating two backdoored models that bypass two state-of-the-art scanners. One model uses callables that are missed by the scanners, and the other model uses disallowed callables by invoking them indirectly. Examples are Hugging Face’s picklescan (Hugging Face, 2024a) and ProtectAI’s scanner (Protect AI, 2024).
@@ 3.1. Study of Pickle Models on Hugging Face
To determine whether pickle models are used despite the availability of alternate formats, we conducted a longitudinal study of the Hugging Face ecosystem. We investigated Hugging Face because it is the largest repository of pre-trained models (Jiang et al., 2022), and because malicious pickled models have been reported on it (Casey et al., 2024; Zhao et al., 2024; Cohen, 2024; Wickens and Bonner, 2024; Zanki, 2025). At 10 points in time over a $\sim$2-year period, we measured the download rates and model formats in repositories with $\geq\!1000$ monthly downloads (as a proxy for real-world impact). [^4] Downloads are tracked using Hugging Face metrics. The number of repositories in a measurement ranged from 2,296 in the first measurement to 16,661 in the last, with the number increasing monotonically at each point. We mined two existing datasets that covered January–October 2023 (PTMTorrent (Jiang et al., 2023b) and HFCommunity (Ait et al., 2023)), and added new measurements in August 2024, November 2024, and March 2025 via the huggingface_hub API. In accordance with previous research (Zhao et al., 2024), we determined model formats using file extensions; interested readers can refer to Appendix˜A for details.
Figure 2. A longitudinal analysis of Hugging Face model formats for repositories with $\geq\!1000$ monthly downloads. Repositories can contain multiple models, each in different formats. Each color groups repositories by the model formats they contain: at least one pickle model (green), exclusively pickle (red), and exclusively SafeTensor (blue).
@@
Figure˜2 summarizes our results. First, as the red lines show, many important models continue to use only the pickle format, and these pickle-only models are downloaded 400M+ times per month. Following the green lines, repositories that contain both pickle and SafeTensor versions of models are also increasingly downloaded, with 1.70 billion monthly downloads. As models are converted to the SafeTensor format, the associated pickle model is often kept for legacy purposes, and can still present security risks (meta llama, 2024).
@@ 3.2.1. Measurements
The weights-only unpickler, introduced in PyTorch 1.13 (Nov. 2022) and enabled by default in PyTorch 2.6 (Nov. 2024) (PyTorch, 2025b), prevents access to insecure callables, but can only load models that use callables from a small allowed set of PyTorch APIs. [^5] PyTorch provides a mechanism for the user to manually expand the set of allowed callables (PyTorch, 2025a), but the user is left to determine by themselves which callables to allow. Models that use more callables cannot be loaded without user intervention; convenience and pressure from end-users results in library maintainers explicitly disabling the weights-only unpickler to maintain compatibility. [^6] As in the case of the flairNLP model (https://github.com/flairNLP/flair/commit/79aa33706e7f753f2edf962feb1d75de22af0d1d). We investigate a sample of the pickle models in our study to determine whether they use callables disallowed by the weights-only unpickler, which affects the usability of the weights-only unpickler as a solution.

### v2

@@
PickleBall: Secure Deserialization of Pickle-based
Machine Learning Models (Extended Report)
Andreas D. Kellas 0009-0001-8763-2400 Columbia UniversityNew YorkNYUSA andreas.kellas@columbia.edu , Neophytos Christou 0000-0001-7335-9485 Brown UniversityProvidenceRIUSA neophytos_christou@brown.edu , Wenxin Jiang 0000-0003-2608-8576 Purdue UniversityWest LafayetteINUSA SocketWilmingtonDEUSA jiang784@purdue.edu , Penghui Li 0000-0002-3077-5697 Columbia UniversityNew YorkNYUSA pl2689@columbia.edu , Laurent Simon 0000-0001-7893-547X GoogleMountain ViewCAUSA laurentsimon@google.com , Yaniv David 0000-0002-1630-7723 TechnionHaifaIsrael yanivmd@technion.ac.il , Vasileios P. Kemerlis 0000-0002-6528-437X Brown UniversityProvidenceRIUSA vpk@cs.brown.edu , James C. Davis 0000-0003-2495-686X Purdue UniversityWest LafayetteINUSA davisjam@purdue.edu and Junfeng Yang 0009-0000-2277-6545 Columbia UniversityNew YorkNYUSA junfeng@cs.columbia.edu
@@ Abstract.
Machine learning model repositories, such as the Hugging Face Model Hub, facilitate model exchanges. However, bad actors can deliver malware through compromised models. Existing defenses, such as safer model formats, restrictive (but inflexible) loading policies, and model scanners, have shortcomings: 44.9% of popular models on Hugging Face still use the insecure pickle format, 15% of these cannot be loaded by restrictive loading policies, and model scanners have both false positives and false negatives. Pickle remains the de facto standard for model exchange, and the ML community lacks a tool that offers transparent safe loading.
We present PickleBall to help machine learning engineers load pickle-based models safely. PickleBall statically analyzes the source code of machine learning libraries and computes custom policies that specify a safe load-time behavior for benign models. It then dynamically enforces these policies during load time as a drop-in replacement for the pickle module. PickleBall generates policies that correctly load 79.8% of benign pickle-based models in our dataset, while rejecting all (100%) malicious examples in the same dataset. In comparison, evaluated model scanners fail to identify known malicious models, and the state-of-the-art loader loads 22% fewer benign models than PickleBall. PickleBall removes the threat of arbitrary function invocation from malicious pickle-based models, raising the bar for attackers as they have to depend on code reuse techniques.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765037isbn: 979-8-4007-1525-9/2025/10ccs: Security and privacy Software and application securityccs: Security and privacy Malware and its mitigation
@@ 1. Introduction
Open-source and open-weight models enable the AI ecosystem (Jiang et al., 2022; Davis et al., 2023; Wang et al., 2022a; Han et al., 2021). They allow machine learning engineers to exchange pre-trained models rather than training from scratch (Dubey et al., 2024), facilitating direct use or fine-tuning (Church et al., 2021). Open model repositories like Hugging Face now host millions of pre-trained models for many tasks (Jiang et al., 2023a, 2022). These model hubs are accessed directly as well as through corporate mirrors (Zhao et al., 2024), with billions of downloads per month.
Similar to traditional software supply-chain attacks, bad actors can distribute malicious models. The most common strategy for achieving remote code execution is tampering with the model deserialization process. Several model serialization formats, such as TorchScript (PyTorch, 2024b), H5/HDF5 (The HDF Group, 2025), and the Python pickle module (Python, 2025b), permit executable metadata or callbacks during model deserialization. Attackers can craft malicious serialized models to execute code, such as os.system(), on victim systems during model loading (Casey et al., 2024; Cohen, 2024; Wickens and Bonner, 2024; Dark Reading Staff, 2024). Researchers have found malicious pickle models on Hugging Face whose payloads include system fingerprinting, credential theft, and reverse shells (Cohen, 2024; Wickens and Bonner, 2024; Zhao et al., 2024), with one study discovering a $5{\times}$ increase in the rate of malicious models uploaded to Hugging Face year-over-year (Zhao et al., 2024).
This led to alternative safe model formats like SafeTensors (Hugging Face, 2024b), and restricted loading APIs like the PyTorch weights-only unpickler (PyTorch, 2024c); we study their adoption in the Hugging Face ecosystem (§3) and find that nevertheless, insecure formats are still prevalent.
In this work, we propose a novel approach to secure pickle model deserialization, which we focus on for three reasons. First, pickle is a popular exchange format for models. Repositories with pickle models are downloaded over 2.1 billion times per month from the Hugging Face model hub (§3.1). Second, pickle is the most expressive format and thus is challenging to secure. Models are encoded as opcodes that are executed by the pickle virtual machine (of Bits, 2024; Huang et al., 2022) during deserialization, which permits invocations of arbitrary Python classes and functions (callables). Third, pickle is abused by attackers. Almost all malicious models on Hugging Face use the pickle format.
@@
To address these limitations, we present PickleBall, a two-part system for securing the exchange of pickle-based models. Our insight is that we can analyze library code to determine the expected behaviors of benign models produced by the library, and enforce tailored model loading policies. PickleBall statically analyzes the library code to learn (1) all class types used in the library’s models and their transitive attribute types, and (2) all functions needed to restore objects of these types. Then, PickleBall’s model-loading component enforces the inferred policies.
We evaluated PickleBall and the state-of-the-art approaches on a dataset of 336 models. Our dataset is meant to represent the kinds of models that PickleBall must handle, including malicious and benign models. We used 252 benign models sourced from Hugging Face, and 84 real and synthetic malicious models. PickleBall loads 79.8% of benign models and prevents all malicious models from executing their payloads. PickleBall adds a runtime overhead of $\sim$2.62% to safely load a model. Compared to other approaches, PickleBall achieved favorable precision and recall.
@@
    The design and implementation of PickleBall, a framework for securely loading pickle models. PickleBall tailors loading policies to models, and enforces these policies lazily, for secure, efficient, and robust model loading. [^1] PickleBall is available at https://github.com/columbia/pickleball.
@@
    A novel dataset of 336 benign and malicious pickle-based models for evaluating pickle security efforts. It has 252 benign models collected from Hugging Face, and 84 malicious models.
@@ 2.1.1. The Model Supply Chain
Training models from scratch requires significant resources (Patterson et al., 2021; Dubey et al., 2024), so engineers and companies reuse machine-learning models trained by others (pre-trained models) (Jiang et al., 2022; Davis et al., 2023; Wang et al., 2022a; Han et al., 2021). Open model repositories like Hugging Face host over 1.8M models (Hugging Face, 2025b) for many tasks (Jiang et al., 2023a, 2022).
A supply chain of pre-trained models has grown from model reuse, which comes with risks similar to those in the traditional software supply chain (Ohm et al., 2020; Jiang et al., 2022). Bad actors apply techniques familiar in traditional software security, such as typosquatting (Neupane et al., 2023; Liu et al., 2022; Jiang et al., 2024, 2025) and code injection (Wang et al., 2022b), as well as ML-specific techniques like model and data manipulation (Gu et al., 2019). Model hubs like Hugging Face try to detect malicious models using both traditional code scanners like ClamAV (Jiang et al., 2022) and domain-specific pickle scanners (AI, 2024a; Hugging Face, 2024a) due to the proliferation of malicious pickle models (Casey et al., 2024; Zhao et al., 2024).
@@ 2.1.2. Model Serialization Formats
Model savers and loaders must agree on the serialization format; there are various formats available, each with its own tradeoffs in terms of security, flexibility, and performance.  [^2] A table showing these tradeoffs is provided in the SafeTensors repository README: https://github.com/huggingface/safetensors. Python is the primary language for using ML models, and its native serialization module, pickle (Python, 2025b), provides a convenient and flexible interface for saving objects; pickle proliferated for being easy to use and is used by popular libraries like PyTorch (PyTorch, 2024a). Hugging Face released the SafeTensors format in September 2022 as an alternative that prioritizes security (Hugging Face, 2024b). The GGUF format, released August 2023, is optimized for fast model loading and inference tasks, especially for large language models (GGML, 2025; Hugging Face, 2025a). Other formats may be selected for library coupling (e.g., TensorFlow SavedModel), interoperability (e.g., ONNX), or intermediate tradeoffs between security, flexibility, and performance.
The security of a format depends on the expressivity of its operations. The SafeTensors format requires very few different operations to load a model, because it only encodes model weight values, and is considered a secure format after independent security audits (Trail of Bits, 2023; Patry, Nicolas and Biderman, Stella, 2023). Some formats, like TensorFlow SavedModel and ONNX, are known to permit operations that could be abused in specific settings (Zhu et al., 2025; Travers, 2021), but with no observed real-world attacks.
Pickle is an extremely expressive format that permits nearly arbitrary operations during deserialization, and numerous malicious pickle models are discovered on Hugging Face (Zhao et al., 2024; Casey et al., 2024; Zanki, 2025; Cohen, 2024; Wickens and Bonner, 2024). We focus our efforts on pickle models due to their insecure format (cf. §2.2) and their continued popularity (cf. §3).
@@ 2.2. Pickle Serialization and Risks
          1  import pickle   2       3  def read_weights_to_tensor(filename: str) -> Tensor:    4  # Read a file containing weights and   5  # return a Tensor object.   6       7  class Tensor(object):    8  ...   9  def __reduce__(self):    10  return (read_weights_to_tensor, (self.filename,))    11       12  class Model(object):    13  def __init__(self, weights: library.Tensor):    14  self.weights = weights    15       16  def save(self, filename):    17  with open(filename, ’wb’) as fd:    18    pickle.dump(self, fd)    19       20  @classmethod   21  def load(cls, path):    22  with open(path, ’rb’) as fd:    23  return pickle.load(fd)    24  ...
@@
Pickle Deserialization Attacks: Pickle’s opcodes allow a pickle program to import and invoke arbitrary Python callables during unpickling. An attacker can use of pickle’s callable importing and callable invoking opcodes to achieve arbitrary code execution—for example by executing an arbitrary shell command by invoking the os.system function. The dangers of deserializing arbitrary pickle programs have been publicized since 2011 (ColdwaterQ, 2022; of Bits, 2024; Elhage, 2023; Python, 2025b).
Protecting Pickle Models: Two existing approaches are used to protect users from malicious pickle models:
@@
    Model scanners identify malicious models using denylists of unsafe callables. As with many denylist approaches, model scanners are useful for identifying recognizable malicious content, but are bypassed by malicious models that avoid denied callables, or that indirectly invoke callables (Tumarkin, 2024). [^3] We demonstrate this by creating two backdoored models that bypass two state-of-the-art scanners. One model uses callables that are missed by the scanners, and the other model uses disallowed callables by invoking them indirectly. See Appendix B. Examples are Hugging Face’s picklescan (Hugging Face, 2024a) and ProtectAI’s scanner (Protect AI, 2024).
@@ 3.1. Study of Pickle Models on Hugging Face
To determine whether pickle models are used despite the availability of alternate formats, we conducted a longitudinal study of the Hugging Face ecosystem. We investigated Hugging Face because it is the largest repository of pre-trained models (Jiang et al., 2022), and because it hosts malicious pickle models (Casey et al., 2024; Zhao et al., 2024; Cohen, 2024; Wickens and Bonner, 2024; Zanki, 2025). At 10 points in time over a $\sim$2-year period, we measured the download rates and model formats in repositories with $\geq\!1000$ monthly downloads (as a proxy for real-world impact). [^4] Downloads are tracked using Hugging Face metrics. The number of repositories in a measurement ranged from 2,296 in the first measurement to 16,661 in the last, with the number increasing monotonically at each point. We mined two existing datasets that covered January–October 2023 (PTMTorrent (Jiang et al., 2023b) and HFCommunity (Ait et al., 2023)), and added new measurements in August 2024, November 2024, and March 2025 via the huggingface_hub API. In accordance with previous research (Zhao et al., 2024), we determined model formats using file extensions; interested readers can refer to Appendix˜A for details.
Figure 2. A longitudinal analysis of Hugging Face model formats for repositories with $\geq\!1000$ monthly downloads. Repositories can contain multiple models, each in different formats. Each color groups repositories by the model formats they contain: at least one pickle model (green), exclusively pickle (red), and exclusively SafeTensors (blue).
@@
Figure˜2 summarizes our results. First, the red lines show that many important models continue to use only the pickle format, and these pickle-only models are downloaded 400M+ times per month. The green lines show that repositories containing both pickle and SafeTensors versions of models are also increasingly downloaded, with 1.70 billion monthly downloads. When models are converted to the SafeTensors format, the associated pickle model is often kept for legacy purposes and can still present security risks (meta llama, 2024).
@@ 3.2.1. Measurements
The weights-only unpickler, introduced in PyTorch 1.13 (Nov. 2022) and enabled by default in PyTorch 2.6 (Nov. 2024) (PyTorch, 2025b), prevents access to dangerous callables, but can only load models that use callables from a small allowed set of PyTorch APIs. [^5] PyTorch provides a mechanism for the user to manually expand the set of allowed callables (PyTorch, 2025a), but the user is left to determine by themselves which callables to allow. Models that use additional callables cannot be loaded without user intervention; convenience and pressure from end-users results in library maintainers explicitly disabling the weights-only unpickler to maintain compatibility. [^6] As in the case of the flairNLP library (https://github.com/flairNLP/flair/commit/79aa33706e7f753f2edf962feb1d75de22af0d1d). We investigate a sample of the pickle models in our study to determine whether they use callables disallowed by the weights-only unpickler, which affects the usability of the weights-only unpickler as a solution.